# Research impact evaluation based on effective authorship contribution sensitivity: h-leadership index

### v1

@@ Abstract
The evaluation of a researcher’s performance has traditionally relied on various bibliometric measures, with the h-index being one of the most prominent. However, the h-index only accounts for the number of citations received in a publication and does not account for other factors such as the number of authors or their specific contributions in collaborative works. Therefore, the h-index has been placed on scrutiny as it has motivated academic integrity issues where non-contributing authors get authorship merely for raising their h-index. In this study, we comprehensively evaluate existing metrics in their ability to account for authorship contribution by their position and introduce a novel variant of the h-index, known as the h-leadership index. The h-leadership index aims to advance the fair evaluation of academic contributions in multi-authored publications by giving importance to authorship position beyond the first and last authors, focused by Stanford’s ranking of the top 2 % of world scientists. We assign weighted citations based on a modified complementary unit Gaussian curve, ensuring that the contributions of middle authors are appropriately recognised. We apply the h-leadership index to analyse the top 50 researchers across the Group of 8 (Go8) universities in Australia, demonstrating its potential to provide a more balanced assessment of research performance. We provide open-source software for extending the work further.
@@ 1 Introduction
The evaluation of a researcher’s performance has used various bibliometric measures based mainly on citations, with the h-index being one of the most prominent. Hirsch introduced the h-index in 2005 [1], which combines a list of publications by a researcher along with the number of citations. The h-index is simple to compute and easy to understand and can be used for evaluating a school, institution, and journal profile, apart from individual researchers. The Google Scholar Metrics  [^1] Google Scholar Metrics: https://scholar.google.com/citations?view_op=top_venues uses the h-index to rank academic journals [2, 3] and provides insights on universities and individual researcher profiles. It has been argued that Google Scholar Metrics can be easily manipulated with poor citation culture and academic dishonesty [4, 5]. Therefore, the h-index has several limitations [6, 7, 8] with the primary criticism being its insensitivity to highly cited papers. Once a paper is included in the h-core (the set of papers that contribute to the h-index), additional citations do not affect the index [9]. This means that even if a paper receives a significant number of citations, the h-index does not reflect this increase, potentially underestimating the impact of influential works. Furthermore, the h-index exhibits a low degree of discrimination, making it less effective in differentiating between researchers with similar citation patterns since authorship contributions (positions) are not taken into account [10]. Another limitation is that the h-index can only increase over time, allowing scientists to rest on their past achievements without accounting for new citations, which may not accurately represent their current research impact [10]; therefore Google Scholar includes the h5-index which looks at citations received in the last five years. Finally, the h-index does not account for the varying contribution of co-authors in multi-authored papers, which can skew the perceived impact of collaborative research efforts [11].
@@
Egghe proposed the g-index in 2006 to address some of the limitations of the h-index by giving more weight to highly cited $g$ papers (g-core) such that together they contribute at least $g^{2}$ citations [12, 13]. However, similar to the h-index, the g-index neglects papers and citations outside the g-core, focusing only on the subset of papers that contribute to the index. This limitation can result in an incomplete assessment of a researcher’s overall impact. Additionally, the g-index can reach a saturation point where subsequent citations do not influence the index unless new papers are published. Jin  [14] introduced the AR-index in 2007, considering the age of publications and giving more weight to recent works that highlight current research activities, it ignores the impact of h-tail papers (papers not in the h-core) and their citations. This can lead to underestimating the research impact for many authors whose influential works lie outside the h-core. Moreover, the AR-index shares some of the h-index’s limitations, such as low discrimination ability and insensitivity to highly cited papers. The AR-index attempts to provide a temporal dimension to citation analysis but still falls short in addressing the complexities of authorship positions and collaborative efforts. This prompted the development of metrics such as the hm-index [11] and h-frac index [15] that utilise weighted citations to provide a more nuanced assessment by considering the number of contributing authors. Recently, more complex measures such as c-score have emerged that provide a composite score combining six factors namely citations, h-index, hm-index, single, first and last authorship percentages [16]. The c-score has gained prominence, ranking top authors by factoring in multiple dimensions of their contributions leading to a more standardised ranking system. An updated version of c-score rankings gets released annually, known as Stanford University’s Top 2% Scientists list [^2] https://elsevier.digitalcommonsdata.com/datasets/btchxktzyw/7. However, despite its robustness, the c-score and similar metrics fall short by inadequately considering the authorship positions beyond the first and last authors. Furthermore, these lists do not apply to authorship culture, where surnames are used to list the authors alphabetically.
@@
The bibliometrics for evaluating academic performance has become increasingly prevalent in recent decades to assess the impact and influence of a university and researcher’s profile. However, the application of bibliometrics involves multiple complexities and loopholes that are easy to exploit, leading to its potential misuse [17]. Therefore, we first explore the development, benefits, and challenges associated with popular bibliometrics, as well as the introduction of composite indicators to provide a more nuanced evaluation of scientific impact. In our study, we introduce a new way to measure research contributions called the h-leadership index. This index assigns credit based on the position of the author such that the first and last authors of a paper get the most credit, the second and second-last authors slightly less, and so on based on a modified complementary unit Gaussian distribution. The h-leadership index assigns weights to an author’s citation for a work based on their authorship position, ensuring the authorship position is accounted for along with the number of citations and number of publications being measured. This new metric thus acknowledges and quantifies the diverse contributions of researchers within collaborative environments. We also present compare the new metrics with more widely used bibliometrics such as h-index [1], hm-index [11] and c-score [16].
@@
As a case study to demonstrate the effectiveness of the h-leadership index, we use Australia’s group of eight (Go8) institutions with retrieved data from the top 50 affiliated researchers. Data was retrieved using Scopus APIs [^3] Elsevier Developer Portal: https://dev.elsevier.com/ which offer generous limits (up to 5000 requests/weekly) for academic research. Go8 universities showcase very high research output with a mean citation of 2.1 million as shown in Figure 1. These institutions are renowned for their substantial contributions to research across various disciplines, making them ideal subjects for examining advanced bibliometric methodologies.
@@ 2 Related Work
The h-index provides a balanced measure by considering both the quantity and the quality of publications. A researcher has a h-index of $h$ if $h$ of their papers have received at least $h$ citations each [18]. Formally, the h-index is the largest number $h$ such that at least $h$ papers have been cited $h$ or more times. It can be expressed mathematically as:
$h=\max_{r}(r\leq c(r))$ (1)
@@
The hm-index, or modified h-index, was proposed as an adjustment to the h-index to address some of its inherent limitations, particularly the equal weighting of all co-authors in multi-authored papers. The hm-index aims to account for the contribution of each co-author by normalising the citation count based on the number of authors. The hm-index is calculated by adjusting the number of citations for each paper according to the number of authors, thus providing a more equitable measure for individual contribution. It is defined as follows: a researcher has a hm-index of $h_{m}$ if $h_{m}$ of their publications have an adjusted citation count (based on the number of authors) that is at least $h_{m}$ [11]. This accounts for varying levels of contribution across papers, particularly in fields where large collaborations are common. It can be expressed mathematically as:
$hm=\max_{r}\left(\sum_{r^{\prime}=1}^{r}\frac{1}{a(r^{\prime})}\leq c(r)\right)$ (2)
@@ 2.2 Composite Indicators
$C_{\text{score}}=\sum_{i=1}^{6}\log\left(\frac{C_{i}}{\max(C_{i})}\right)$ (3)
@@ 3.1 Framework for the h-leadership index
, where the peak is at the ends, rather than in the middle. This transformation effectively inverts the bell curve, giving higher values to the tails (i.e., positions at the start and end). The lowest value in complementary gaussian distribution is 0 which can leave some authors with no contribution. Thus, we modified complementary unit Gaussian distribution as:
@@
Figure 2: Variation of weights assigned to authors based on their authorship position in the h-leadership index. First and last authorship position get the highest contribution at 1 which marginal decrement as the authorship position moves towards the middle authors. The leadership weight assigned at position 5 is 0.992 which reaches 0.3 by position 50. If the position from first and last authors is greater than 50, then the weight stays static at 0.3 (lowest assigned).
The choice of a modified complementary unit Gaussian curve for weighting is justified as it naturally emphasises the first and last authors, who typically lead and finalise the research while reducing the weight for middle authors who may have had more supportive roles. Other potential models, such as linear or exponential decay functions, were considered but found to either overly penalise or under represent certain authorship positions, leading to less accurate assessments. The modified complementary unit Gaussian model offers the best balance, accurately reflecting the typical contributions across different authorship positions.
Using the modified complementary unit Gaussian distribution, we calculate the h-leadership index by first assigning weights to citations based on the author’s position in the author list. We then soft the list of weighted citations in descending order, and then determine the h-leadership index as the largest number $h$ such that the author has $h$ papers, each with at least $h$ weighted citations. Mathematically, let $C_{i}$ be the citations received by the $i$-th paper, and $w(x)$ be the weight assigned based on the author’s position $x$. The h-leadership index $h_{L}$ is then given by:
@@
Algorithm 1 calculates the h-leadership index for a given author, identified by their scopus_id, based on a list of their publications. The algorithm works by iterating through each publication to compute weighted citations for the author. It then computes the weight based on the author’s position in the list of co-authors, using a function leadership_weight(x: author position, n: number of authors in a publication). The citations are accumulated and sorted in descending order and the h-leadership index is determined as the maximum number of publications with at least $h$ citations, adjusted by the author’s position in the author list.
@@ 3.2 Data retrieval
We retrieved the data for this using Scopus, a comprehensive database for academic research outputs as shown in Algorithm 2. The focus was on retrieving publication records for the top 50 researchers based on citations affiliated with Go8. We identified the top researchers at these universities using their Google Scholar affiliation page, and a list of researchers from these universities. The need for using Google Scholar in determining the top researchers arose from the observation that Scopus sometimes list scholars in the same university as affiliated with their school instead of the university making it difficult to group them. Our approach ensures a robust dataset encompassing a diverse range of publication records and accurate citation metrics essential for computing the proposed h-leadership index. Using the Scopus API, we searched for authors based on their names, affiliations, and countries. For each identified author, we extracted publication meta-data, including titles, authorship details, citation counts, publication venues, and other relevant metadata. The retrieved data was stored in a structured format, enabling subsequent analysis. We also compiled the list of the top domestic and international collaborators for these researchers to analyse the overall networking that pertains to modern research studies. The output of this process is a list of correctly matched researchers for each affiliation, along with their complete publication history, obtained through iterative requests from the Scopus API as shown in Figure 3. We selected Scopus as the primary data source due to its extensive coverage of scholarly literature and reliable citation-tracking capabilities. Despite challenges related to affiliation consistency, Scopus remains integral to bibliometric research, offering access to a vast database crucial for conducting comprehensive analyses of publication impact and authorship dynamics [27].
@@
We retrieved a total of 400 researcher records (50 each from the eight institutions) with information on their Scopus ID (identifier), name, publications, affiliation, city, country, number of documents published, and subject areas. Each publication includes information on the article title, co-authors, citations, publication name, venue, cover date, ISSN (International Standard Serial Number), issue, page range and DOI (Digital Object Identifier). The co-author’s affiliations are also listed.
@@ 3.3 Technical details
The experimental process shown in Figure 4 consists of three main stages: managing researcher affiliations, retrieving publication data from the Scopus API, and calculating bibliometric indicators. Our setup is flexible, allowing for adaptation to different sets of researchers and institutions and provided in the code implementation on our GitHub repository [^4] https://github.com/nepython/metrics.git. We store the mapping between universities and their associated researchers in a configuration file, affiliations.json, located in the data directory. Each university is a key, with its corresponding researchers listed as values and to adapt the data extraction for a different set of universities and researchers, one must modify this file by adding the relevant institution names and updating the associated list of researchers. An example structure of the affiliations.json file is as follows:
@@ 4.1 Data and bibliometrics
We first analyse the publication data of top-cited 400 Go8 Australian university researchers (50 per university), encompassing a total of 168,563 scholarly works. Among these, 154,337 (91.56%) were published in journals, 8,517 (5.05%) in conference proceedings, and 5,471 (3.25%) as books or book series. Other venues included reports (0.004%), trade journals (0.09%), and unknown sources (0.05%).
@@
Table 1 presents a comparative summary of bibliometric metrics computed using this publication data. We observe a strong correlation between h-leadership and h-index with the h-index decrementing slightly with the increase in the number of overall authors and a decrease in authorship contribution as measured by the authorship position. Notably, the h-leadership index consistently provides more granular differentiation compared to the h-index, particularly for researchers with high collaborative outputs. For instance, UNSW has the lowest median number of co-authors amongst papers analysed which is also reflected in their h-leadership index (77) being close to h-index (80) post weighted reduction. That being said, the number of co-authors isn’t the sole criterion determining individual contribution. This is reflected in UniMelb’s h-leadership index (74) remaining high compared to the h-index (78), despite its researchers having 11 median co-authors per paper. Comparatively, USyd having a similar number of median co-authors per paper (12), scores 74 in the h-leadership index against its h-index of 86. This emphasises the fairer assessment enabled by the weighted citation scheme which strikes a balance between the number of co-authors and the authorship position, not adversely penalising larger study groups.
@@
Table 2 presents the hl-index for leading researchers defined by citations received on researchers’ Google Scholar profiles in the selected Go8 Australian universities. For instance, Google Scholar provides a list of the top 10 researchers at UniMelb based on citations with links on the page to the next 10 researchers  [^5] https://scholar.google.com/citations?view_op=view_org&hl=en&org=10194437184893062620. The reason for using Google Scholar for top author identification instead of relying on Scopus has been explained in Section 3.2. We selected 10 out of the top 50 profiles in this study based on random sampling from each of the Go8 Australian universities in Table 2. We note that the top-cited researchers at Australian National University (ANU) and Monash University engage in Physics, Astronomy and Medicine disciplines (Figure 9), which explains the higher number of co-authors observed. We also note that Scopus search API clips a publication’s authors to the first 100 and obtaining a full list requires using abstract retrieval API [^6] https://cran.r-project.org/web/packages/rscopus/vignettes/multi_author.html. Since abstract retrieval API call needs to be made per publication, we decided against its usage due to existing weekly quota limits. We highlight that in our selected random samples, we have 2 author profiles with 100 median coauthors at ANU and 3 at Monash. We notice huge variations when comparing the h-index to our hl-index (59 vs 4) for the 3rd ANU profile and the 5th ANU profile (80 vs 5). We also notice similar patterns when comparing the h-index and hl-index for Monash profiles (1st, 3rd, and 5th profiles. We provide a complete list of the top 50 profiles in metrics.csv for all Go8 universities via our GitHub repository  [^7] https://github.com/nepython/metrics/tree/main/results. We also notice that the first, single and last authorship makes a large impact on the hl-index when compared to the h-index, since authorship position is the key principle of the hl-index. In case when medium number of coauthors is low (2 or 3), then the hl-index is close to the h-index.
@@
Table 2: h-leadership index for 10 randomly sampled researchers across a) UniMelb, b) USyd, c) ANU, d) Monash and e) UNSW. We have not listed any metrics such as citations, or number of publications to protect author privacy. We derive the median citation number by sorting an author’s publications by citation and considering the median publication’s citation count. We derive the median author position by sorting publications based on authorship position. We sort publications by the number of coauthors to obtain the median number of coauthors.

### v2

@@ Abstract
The evaluation of a researcher’s performance has traditionally relied on various bibliometric measures, with the h-index being one of the most prominent. However, the h-index only accounts for the number of citations received in a publication and does not account for other factors, such as the number of authors or their specific contributions in collaborative works. Therefore, the h-index has been placed under scrutiny as it has motivated academic integrity issues where non-contributing authors get authorship merely for raising their h-index. In this study, we comprehensively evaluate existing metrics’ ability to account for authorship contribution by their position and introduce a novel h-index variant, the h-leadership index. The h-leadership index aims to advance the fair evaluation of academic contributions in multi-authored publications by giving importance to authorship position beyond the first and last authors, focused by Stanford’s ranking of the top 2 % of world scientists. We assign weighted citations based on a modified complementary unit Gaussian curve, ensuring that the contributions of middle authors are appropriately recognised. We apply the h-leadership index to analyse the top 50 researchers across the Group of 8 (Go8) universities in Australia, demonstrating its potential to provide a more balanced assessment of research performance. We provide open-source software to extend the work further.
@@ 1 Introduction
The evaluation of a researcher’s performance has used various bibliometric measures based mainly on citations, with the h-index being one of the most prominent. Hirsch introduced the h-index in 2005 [1], which combines a list of publications by a researcher along with the number of citations. The h-index is simple to compute and easy to understand and can be used for evaluating a school, institution, and journal profile, apart from individual researchers. The Google Scholar Metrics  [^1] Google Scholar Metrics: https://scholar.google.com/citations?view_op=top_venues uses the h-index to rank academic journals [2, 3] and provides insights on universities and individual researcher profiles. It has been argued that Google Scholar Metrics can be easily manipulated with poor citation culture and academic dishonesty [4, 5]. Therefore, the h-index has several limitations [6, 7, 8] with the primary criticism being its insensitivity to highly cited papers. Once a paper is included in the h-core (the set of papers that contribute to the h-index), additional citations do not affect the index [9]. This means that even if a paper receives a significant number of citations, the h-index does not reflect this increase, potentially underestimating the impact of influential works. Furthermore, the h-index exhibits a low degree of discrimination, making it less effective in differentiating between researchers with similar citation patterns since authorship contributions (positions) are not taken into account [10]. Another limitation is that the h-index can only increase over time, allowing scientists to rest on their past achievements without accounting for new citations, which may not accurately represent their current research impact [10]; therefore, Google Scholar includes the h5-index, which looks at citations received in the last five years. Finally, the h-index does not account for the varying contribution of co-authors in multi-authored papers, which can skew the perceived impact of collaborative research efforts [11].
@@
Egghe proposed the g-index in 2006 to address some of the limitations of the h-index by giving more weight to highly cited $g$ papers (g-core) such that together they contribute at least $g^{2}$ citations [12, 13]. However, similar to the h-index, the g-index neglects papers and citations outside the g-core, focusing only on the subset of papers that contribute to the index. This limitation can result in an incomplete assessment of a researcher’s overall impact. Additionally, the g-index can reach a saturation point where subsequent citations do not influence the index unless new papers are published. Jin  [14] introduced the AR-index in 2007, considering the age of publications and giving more weight to recent works that highlight current research activities. It ignores the impact of h-tail papers (papers not in the h-core) and their citations. This can lead to underestimating the research impact for many authors whose influential works lie outside the h-core. Moreover, the AR index shares some of the h-index’s limitations, such as low discrimination ability and insensitivity to highly cited papers. The AR-index attempts to provide a temporal dimension to citation analysis but still falls short in addressing the complexities of authorship positions and collaborative efforts. This prompted the development of metrics such as the hm-index [11] and h-frac index [15] that utilise weighted citations to provide a more nuanced assessment by considering the number of contributing authors. Recently, more complex measures such as c-score have emerged that provide a composite score combining six factors, namely citations, h-index, hm-index, single, first and last authorship percentages [16]. The c-score has gained prominence, ranking top authors by factoring in multiple dimensions of their contributions, leading to a more standardised ranking system. An updated version of c-score rankings gets released annually, known as Stanford University’s Top 2% Scientists list [^2] https://elsevier.digitalcommonsdata.com/datasets/btchxktzyw/7. However, despite its robustness, the c-score and similar metrics fall short by inadequately considering the authorship positions beyond the first and last authors. Furthermore, these lists do not apply to authorship culture, where surnames are used to list the authors alphabetically.
@@
The bibliometrics for evaluating academic performance has become increasingly prevalent in recent decades to assess the impact and influence of a university and researcher’s profile. However, the application of bibliometrics involves multiple complexities and loopholes that are easy to exploit, leading to its potential misuse [17]. Therefore, we first explore the development, benefits, and challenges associated with popular bibliometrics, as well as the introduction of composite indicators to provide a more nuanced evaluation of scientific impact. In our study, we introduce a new way to measure research contributions called the h-leadership index. This index assigns credit based on the position of the author such that the first and last authors of a paper get the most credit, the second and second-last authors slightly less, and so on, based on a modified complementary unit Gaussian distribution. The h-leadership index assigns weights to an author’s citation for a work based on their authorship position, ensuring the authorship position is accounted for along with the number of citations and number of publications being measured. This new metric thus acknowledges and quantifies the diverse contributions of researchers within collaborative environments. We also present and compare the new metrics with more widely used bibliometrics such as h-index [1], hm-index [11] and c-score [16].
@@
As a case study to demonstrate the effectiveness of the h-leadership index, we use Australia’s group of eight (Go8) institutions with retrieved data from the top 50 affiliated researchers. Data was retrieved using Scopus APIs [^3] Elsevier Developer Portal: https://dev.elsevier.com/, which offer generous limits (up to 5000 requests/weekly) for academic research. Go8 universities showcase very high research output with a mean citation of 2.1 million as shown in Figure 1. These institutions are renowned for their substantial contributions to research across various disciplines, making them ideal subjects for examining advanced bibliometric methodologies.
@@ 2 Related Work
The h-index provides a balanced measure by considering both the quantity and the quality of publications. A researcher has an h-index of $h$ if $h$ of their papers have received at least $h$ citations each [18]. Formally, the h-index is the largest number $h$ such that at least $h$ papers have been cited $h$ or more times. It can be expressed mathematically as:
$h=\max_{r}(r\mid r\leq c(r))$ (1)
@@
The hm-index, or modified h-index, was proposed as an adjustment to the h-index to address some of its inherent limitations, particularly the equal weighting of all co-authors in multi-authored papers. The hm-index aims to account for the contribution of each co-author by normalising the citation count based on the number of authors. The hm-index is calculated by adjusting the number of citations for each paper according to the number of authors, thus providing a more equitable measure for individual contribution. It is defined as follows: a researcher has an hm-index of $h_{m}$ if $h_{m}$ of their publications have an adjusted citation count (based on the number of authors) that is at least $h_{m}$ [11]. This accounts for varying levels of contribution across papers, particularly in fields where large collaborations are common. It can be expressed mathematically as:
$h_{m}=\max_{r}\left(r\mid\sum_{r^{\prime}=1}^{r}\frac{1}{a(r^{\prime})}\leq c(r)\right)$ (2)
@@ 2.2 Composite Indicators
$C_{\text{score}}=\sum_{i=1}^{6}\frac{\log(1+C_{i})}{\max\log(1+C_{i})}$ (3)
@@ 3.1 Framework for the h-leadership index
, where the peak is at the ends rather than in the middle. This transformation effectively inverts the bell curve, giving higher values to the tails (i.e., positions at the start and end). The lowest value in complementary Gaussian distribution is 0, which can leave some authors with no contribution. Thus, we modified the complementary unit Gaussian distribution as:
@@
Figure 2: Variation of weights assigned to authors based on their authorship position in the h-leadership index. First and last authorship positions get the highest contribution at 1, with a marginal decrement as the authorship position moves towards the middle authors. The leadership weight assigned at position 5 is 0.992, which reaches 0.3 by position 50. If the position from the first and last authors is greater than 50, then the weight stays static at 0.3 (lowest assigned).
The choice of a modified complementary unit Gaussian curve for weighting is justified as it naturally emphasises the first and last authors, who typically lead and finalise the research while reducing the weight for middle authors who may have had more supportive roles. Other potential models, such as linear or exponential decay functions, were considered but found to either overly penalise or under-represent certain authorship positions, leading to less accurate assessments. The modified complementary unit Gaussian model offers the best balance, accurately reflecting the typical contributions across different authorship positions.
Using the modified complementary unit Gaussian distribution, we calculate the h-leadership index by first assigning weights to citations based on the author’s position in the author list. We then sort the list of weighted citations in descending order and then determine the h-leadership index as the largest number $h$ such that the author has $h$ papers, each with at least $h$ weighted citations. Mathematically, let $C_{i}$ be the citations received by the $i$-th paper, and $w(x)$ be the weight assigned based on the author’s position $x$. The h-leadership index $h_{L}$ is then given by:
@@
Algorithm 1 calculates the h-leadership index for a given author, identified by their scopus_id, based on a list of their publications. The algorithm works by iterating through each publication to compute weighted citations for the author. It then computes the weight based on the author’s position in the list of co-authors, using a function leadership_weight(x: author position, n: number of authors in a publication). The citations are accumulated and sorted in descending order, and the h-leadership index is determined as the maximum number of publications with at least $h$ citations, adjusted by the author’s position in the author list.
@@ 3.2 Data retrieval
We retrieved the data for this using Scopus, a comprehensive database for academic research outputs, as shown in Algorithm 2. The focus was on retrieving publication records for the top 50 researchers based on citations affiliated with Go8. We identified the top researchers at these universities using their Google Scholar affiliation page and a list of researchers from these universities. The need for using Google Scholar in determining the top researchers arose from the observation that Scopus sometimes list scholars in the same university as affiliated with their school instead of the university, making it difficult to group them. Our approach ensures a robust dataset encompassing a diverse range of publication records and accurate citation metrics essential for computing the proposed h-leadership index. Using the Scopus API, we searched for authors based on their names, affiliations, and countries. For each identified author, we extracted publication meta-data, including titles, authorship details, citation counts, publication venues, and other relevant metadata. The retrieved data was stored in a structured format, enabling subsequent analysis. We also compiled the list of the top domestic and international collaborators for these researchers to analyse the overall networking that pertains to modern research studies. The output of this process is a list of correctly matched researchers for each affiliation, along with their complete publication history, obtained through iterative requests from the Scopus API as shown in Figure 3. We selected Scopus as the primary data source due to its extensive coverage of scholarly literature and reliable citation-tracking capabilities. Despite challenges related to affiliation consistency, Scopus remains integral to bibliometric research, offering access to a vast database crucial for conducting comprehensive analyses of publication impact and authorship dynamics [27].
@@
We retrieved a total of 400 researcher records (50 each from the eight institutions) with information on their Scopus ID (identifier), name, publications, affiliation, city, country, number of documents published, and subject areas. Each publication includes information on the article title, co-authors, citations, publication name, venue, cover date, ISSN (International Standard Serial Number), issue, page range and DOI (Digital Object Identifier). The co-authors’ affiliations are also listed.
@@ 3.3 Technical details
The experimental process shown in Figure 4 consists of three main stages: managing researcher affiliations, retrieving publication data from the Scopus API, and calculating bibliometric indicators. Our setup is flexible, allowing for adaptation to different sets of researchers and institutions and provided in the code implementation on our GitHub repository [^4] https://github.com/nepython/metrics.git. We store the mapping between universities and their associated researchers in a configuration file, affiliations.json, located in the data directory. Each university is a key, with its corresponding researchers listed as values, and to adapt the data extraction for a different set of universities and researchers, one must modify this file by adding the relevant institution names and updating the associated list of researchers. An example structure of the affiliations.json file is as follows:
@@ 4.1 Data and bibliometrics
We first analyse the publication data of the top-cited 400 Go8 Australian university researchers (50 per university), encompassing a total of 168,563 scholarly works. Among these, 154,337 (91.56%) were published in journals, 8,517 (5.05%) in conference proceedings, and 5,471 (3.25%) as books or book series. Other venues included reports (0.004%), trade journals (0.09%), and unknown sources (0.05%).
@@
Table 1 presents a comparative summary of bibliometric metrics computed using this publication data. We observe a strong correlation between h-leadership and h-index, with the h-index decrementing slightly with the increase in the number of overall authors and a decrease in authorship contribution as measured by the authorship position. Notably, the h-leadership index consistently provides more granular differentiation compared to the h-index, particularly for researchers with high collaborative outputs. For instance, UNSW has the lowest median number of co-authors amongst papers analysed, which is also reflected in their h-leadership index (77) being close to h-index (80) post-weighted reduction. That being said, the number of co-authors isn’t the sole criterion determining individual contribution. This is reflected in UniMelb’s h-leadership index (74) remaining high compared to the h-index (78), despite its researchers having 11 median co-authors per paper. Comparatively, USyd, having a similar number of median co-authors per paper (12), scores 74 in the h-leadership index against its h-index of 86. This emphasises the fairer assessment enabled by the weighted citation scheme, which strikes a balance between the number of co-authors and the authorship position, not adversely penalising larger study groups.
@@
Table 2 presents the hl-index for leading researchers defined by citations received on researchers’ Google Scholar profiles in the selected Go8 Australian universities. For instance, Google Scholar provides a list of the top 10 researchers at UniMelb based on citations with links on the page to the next 10 researchers  [^5] https://scholar.google.com/citations?view_op=view_org&hl=en&org=10194437184893062620. The reason for using Google Scholar for top author identification instead of relying on Scopus has been explained in Section 3.2. We selected 10 out of the top 50 profiles in this study based on random sampling from each of the Go8 Australian universities in Table 2. We note that the top-cited researchers at Australian National University (ANU) and Monash University engage in Physics, Astronomy and Medicine disciplines (Figure 9), which explains the higher number of co-authors observed. We also note that Scopus search API clips a publication’s authors to the first 100, and obtaining a full list requires using abstract retrieval API [^6] https://cran.r-project.org/web/packages/rscopus/vignettes/multi_author.html. Since an abstract retrieval API call needs to be made per publication, we decided against its usage due to existing weekly quota limits. We highlight that in our selected random samples, we have 2 author profiles with 100 median coauthors at ANU and 3 at Monash. We notice huge variations when comparing the h-index to our hl-index (59 vs 4) for the 3rd ANU profile and the 5th ANU profile (80 vs 5). We also notice similar patterns when comparing the h-index and hl-index for Monash profiles (1st, 3rd, and 5th profiles. We provide a complete list of the top 50 profiles in metrics.csv for all Go8 universities via our GitHub repository  [^7] https://github.com/nepython/metrics/tree/main/results. We also notice that the first, single and last authorship makes a large impact on the hl-index when compared to the h-index, since authorship position is the key principle of the hl-index. In cases when a medium number of coauthors is low (2 or 3), then the hl-index is close to the h-index.
@@
Table 2: h-leadership index for 10 randomly sampled researchers across a) UniMelb, b) USyd, c) ANU, d) Monash and e) UNSW. We have not listed any metrics such as citations or number of publications to protect author privacy. We derive the median citation number by sorting an author’s publications by citation and considering the median publication’s citation count. We derive the median author position by sorting publications based on authorship position. We sorted publications by the number of coauthors to obtain the median number of coauthors.